# On the generalized circular projected Cauchy distribution

### 10 Mar 2026 v2

@@ Abstract
Tsagris and Alzeley, (2025) proposed the generalized circular projected Cauchy distribution, whose special case is the wrapped Cauchy distribution. In this paper we first derive the relationship with the wrapped Cauchy distribution, and then we attempt to characerize the distribution. We provide a non-analytical formula for the mean resultant length, and analytical form for the cumulative probability function and the the entropy. We propose a log-likelihood ratio test for the equality of two angular means, without assuming equality of thew concentration parameters, and briefly mention the addition of circular and simplicial predictors. Simulation studies illustrate a) the performance of the log-likelihood ratio test when one falsely assumes that the true distribution is the wrapped Cauchy distribution, and b) the empirical rate of covergence of the regression coefficients. Using a real data analysis example we correct a mistake in the regression setting.
Keywords: Circular data; mean resultant length, hypothesis test
@@ 1 Introduction
where $\left|\left|.\right|\right|$ denotes the Euclidean norm. Circular data, when $d=2$, lie on a circle. Circular data are encountered in various disciplines, such as political sciences (Gill and Hangartner,, 2010), criminology (Shirota and Gelfand,, 2017), biology (Landler et al.,, 2018), ecology (Horne et al.,, 2007), and astronomy (Soler et al.,, 2019), to name but a few.
A large class of distributions has been proposed, see Tsagris and Alzeley, (2025) for a short list. In this paper we focus on the generalized circular projected Cauchy distribution (GCPC) proposed by Tsagris and Alzeley, (2025). Specifically we show the relationship with the wrapped Cauchy (WC) distribution and further propose a log-likelihood ratio test for the equality of two independent angular means.
The next section briefly presents the GCPC distribution, derives the relationship with the WC distribution and proposes the log-likelihood ratio test. Then, simulation studies follow and the conclusions that close the paper.
@@ 2 The GCPC distribution
By substituting (2) into (1) and evaluating the integral, Tsagris and Alzeley, (2025) derived the termed the Circular Projected Cauchy (CPC) distribution.
@@
(Tsagris and Alzeley,, 2025) relaxed this strict assumption by employing one of the conditions imposed in Paine et al., (2018), that is, $\boldsymbol{\Sigma\mu}=\boldsymbol{\mu}$, but not $|\boldsymbol{\Sigma}|=1$. This condition implies that the one eigenvector $\boldsymbol{\xi}_{2}$ of $\boldsymbol{\Sigma}$ is the normalised location vector $\boldsymbol{\xi}_{2}=\left(\mu_{1},\mu_{2}\right)^{\top}/\gamma$, while the other eigenvector can be defined up to the sign as $\boldsymbol{\xi}_{1}=\left(-\mu_{2},\mu_{1}\right)^{\top}/\gamma$ or $\boldsymbol{\xi}_{1}=\left(\mu_{2},-\mu_{1}\right)^{\top}/\gamma$. The eigenvalue corresponding to the location vector is equal to 1, while the other eigenvalue is equal to $\lambda$, hence $|\boldsymbol{\Sigma}|=\lambda>0$ and the inverse of the scatter matrix is given by
@@
The GCPC distribution exhibits symmetry with respect to $\omega$ such that $f(\theta-\omega)=f(\omega-\theta)$, and is thus symmetrical about $\omega+\pi$. The density of the GCPC may be re-written as
@@
It is important to note that if $\lambda=1$, the GCPC distribution reduces to the circular independent projected Cauchy (CIPC) distribution (Tsagris et al.,, 2025)
@@
which is the WC distribution (Mardia and Jupp,, 2000, pg. 51) with a different parameterization
@@
where $\gamma=\frac{2\delta}{1-\delta^{2}}$ or, conversely, $\delta=(\sqrt{\gamma^{2}+1}-1)/\gamma$ (Tsagris and Alzeley,, 2025).
@@ 2.1 Relationship with the CIPC distribution
The proof is straightfoward by application of the change-of-variables formula to obtain
$f_{\Psi}(\psi)=f_{\Theta}(\theta)\cdot\left|\frac{d\theta}{d\psi}\right|=\frac{1}{2\pi\sqrt{\lambda}\,\sqrt{b}\,D}\cdot\sqrt{\lambda}\cdot b=\frac{1}{2\pi\!\left(\sqrt{\gamma^{2}+1}-\gamma\cos\psi\right)}.$
@@ 2.2 Probabilities
Following Tsagris and Alzeley, (2025), instead of the cumulative probability function we derive the probability included within a given interval $(a,b)$, and provide an alternative formula.
@@ 2.3 Mean resultant length
The mean resultant length is defined as $\rho=E\left[\cos{\left(\theta-\omega\right)}\right]$. Based on this, one may compute the circular variance as $1-\rho$, and the circular standard deviation as $\left(-2\log{\rho}\right)^{1/2}$. Using Theorem 2.1 we may express this expectation as
$\rho=\frac{1-\delta^{2}}{\pi}\int_{0}^{\pi}\frac{\cos\psi}{\sqrt{\cos^{2}\psi+\lambda\sin^{2}\psi}}\cdot\frac{1}{1+\delta^{2}-2\delta\cos\psi}\,d\psi.$
After some algebra we end up with the following representation
$\rho=\frac{2(1+\delta)}{\pi(1-\delta)\sqrt{\lambda}}\;\Pi\!\left(\left(\frac{1+\delta}{1-\delta}\right)^{2}\frac{1}{\lambda}-1\;\Bigg|\;1-\frac{1}{\lambda^{2}}\right),$
where $\Pi(a|b)$ is the complete elliptic integral of the third kind (Ward,, 1960). We still see that there is no closed form solution for $\rho$, unless $\lambda=1$, and this applies to higher trigonometric moments. Figure 1(a) visualizes the values of $\rho$ for a grid of values of $\gamma$ and $\lambda$. We observe that $\rho$ increases with increasing $\gamma$ and decreases with increasing $\lambda$ values.
@@ 2.4 Entropy
$H_{GCPC}(\omega,\gamma,\lambda)=\log{\left\{\frac{8\pi\sqrt{\lambda}\,(1-\delta^{2})}{\bigl[\sqrt{\lambda}(1-\delta^{2})+1+\delta^{2}\bigr]^{2}}\right\}},$
where $\delta$ was defined earlier.
@@
Note that if $\lambda=1$: $H_{GCPC}=\log(2\pi(1-\delta^{2}))=H_{WC}$, and if $\delta=0$: $H_{GCPC}=\log(8\pi\sqrt{\lambda}/(\sqrt{\lambda}+1)^{2})\leq\log(2\pi)$. Figure 1(b) visualizes the values of $H$ for a grid of values of $\gamma$ and $\lambda$. We observe that, in cotrast to $\rho$, $H$ increases with increasing $\lambda$ and decreases with increasing $\gamma$ values.
@@ 2.6.1 Hypothesis test for the location parameter of one sample
In order to test whether the location parameter equals some pre-specified value we will employ the log-likelihood ratio test. Under the $H_{0}$, $\omega=\omega_{0}$, and so we need to maximize the restricted log-likelihood, $\ell_{0}$, with respect to the other two parameters, $\gamma$ and $\lambda$, and estimate their values, $\tilde{gamma}$ and $\tilde{\lambda}$. Under the $H_{1}$, $\omega\neq\omega_{0}$, and we maximize the unrestricted log-likelihood, $\ell_{1}$, to obtain $\hat{\omega}$, $\hat{\gamma}$ and $\hat{\lambda}$. Under regularity conditions, $\Lambda=2\left[\ell_{1}\left(\hat{\omega},\hat{\omega},\hat{\lambda}\right)-\ell_{0}\left(\omega_{0},\tilde{\gamma},\tilde{\lambda}\right)\right]\mathrel{\dot{\sim}}\chi^{2}_{1}$.
2.6.2 Confidence interval for a location parameter
Using the log-likelihood ratio test, we can construct asymptotic confidence intervals for the true location parameter $\omega$ of a sample by searching for the pair of values for which it holds that (Cox and Hinkley,, 1979): $\ell\left(\omega,\hat{\gamma},\hat{\lambda}\right)>\ell\left(\hat{\omega},\hat{\gamma},\hat{\lambda}\right)-\frac{\chi^{2}_{1}}{2}$.

### 25 Mar 2026 v3

@@ Abstract
Tsagris and Alzeley (2025) proposed the generalized circular projected Cauchy (GCPC) distribution, whose special case is the wrapped Cauchy distribution. In this paper we first derive the relationship with the wrapped Cauchy distribution, and then we attempt to characterize the distribution. We establish the conditions under which the distribution exhibits unimodality. We provide non-analytical formulas for the mean resultant length and the Kullback-Leibler divergence, and analytical form for the cumulative probability function and the entropy of the GCPC distribution. We propose log-likelihood ratio tests for one, or two location parameters without assuming equality of the concentration parameters. We revisit maximum likelihood estimation with and without predictors. In the regression setting we briefly mention the addition of circular and simplicial predictors. Simulation studies illustrate a) the performance of the log-likelihood ratio test when one falsely assumes that the true distribution is the wrapped Cauchy distribution, and b) the empirical rate of convergence of the regression coefficients. Using a real data analysis example we show how to avoid the log-likelihood being trapped in a local maximum and we correct a mistake in the regression setting.
Keywords: Circular data; distribution, hypothesis test, maximum likelihood estimation
@@ 1 Introduction
where $\left|\left|.\right|\right|$ denotes the Euclidean norm. Circular data, when $d=2$, lie on a circle. Circular data are encountered in various disciplines, such as political sciences (Gill and Hangartner, 2010), criminology (Shirota and Gelfand, 2017), biology (Landler et al., 2018), ecology (Horne et al., 2007), and astronomy (Soler et al., 2019), to name but a few.
A large class of distributions has been proposed, see Tsagris and Alzeley (2025) for a short list. A classic distribution is the wrapped Cauchy distribution (Kent and Tyler, 1988), for which Tsagris and Alzeley (2025) showed to be a special case of the generalized projected Cauchy distribution (GCPC). The GCPC generalizes the WC by the introduction of an extra parameter that allows for anisotropy. The benefit of the GCPC distribution is that it provides a better fit to asymmetric and bimodal data.
In this paper we focus on the GCPC distribution. Specifically we derive the relationship with the wrapped Cauchy (WC) distribution. We examine the conditions for unimodality and then provide an alternative formula for the cumulative probability function. We derive non-analytical formulas for its mean resultant length and the Kullback-Leibler divergence from the WC distribution, but derive analytical formula for its entropy. We further propose two log-likelihood ratio tests for equality of one and two location parameters, without assuming equal concentration parameters. We revisit the maximum likelihood estimation (MLE) and the regression modelling. For the MLE we show a problem that can trap the log-likelihood in a local maximum and show how to easily escape and reach the global maximum. We further correct a mistake of Tsagris and Alzeley (2025) and examine empirically the convergence rate of the regression coefficients. A real data analysis, with and without predictor variables illustrates the superior performance of the GCPC compared to the WC distribution.
The next section briefly presents the GCPC distribution, and studies the distribution. Next, simulation studies and the real data example follow, with conclusions closing the paper.
@@ 2 The GCPC distribution
By substituting (2) into (1) and evaluating the integral, Tsagris and Alzeley (2025) derived the circular projected Cauchy (CPC) distribution.
@@
(Tsagris and Alzeley, 2025) relaxed this strict assumption by employing one of the conditions imposed in Paine et al. (2018), that is, $\boldsymbol{\Sigma\mu}=\boldsymbol{\mu}$, but not $|\boldsymbol{\Sigma}|=1$. This condition implies that the one eigenvector $\boldsymbol{\xi}_{2}$ of $\boldsymbol{\Sigma}$ is the normalised location vector $\boldsymbol{\xi}_{2}=\left(\mu_{1},\mu_{2}\right)^{\top}/\gamma$, while the other eigenvector can be defined up to the sign as $\boldsymbol{\xi}_{1}=\left(-\mu_{2},\mu_{1}\right)^{\top}/\gamma$ or $\boldsymbol{\xi}_{1}=\left(\mu_{2},-\mu_{1}\right)^{\top}/\gamma$. The eigenvalue corresponding to the location vector is equal to 1, while the other eigenvalue is equal to $\lambda$, hence $|\boldsymbol{\Sigma}|=\lambda>0$ and the inverse of the scatter matrix is given by
@@
We shall denote the eigenvalue, $\lambda$, of the covariance matrix $\bm{\Sigma}$ by anisotropy parameter, and the reason is explained in the next subsection. The GCPC distribution exhibits reflective symmetry with respect to $\omega$ only if $\lambda=1$, but its density function is even since $f(\theta-\omega)=f(\omega-\theta)$. The maximum value of the density occurs when $\theta=\omega$ and its value is $\left[2\pi\sqrt{\lambda}\left(\sqrt{\gamma^{2}+1}-\gamma\right)\right]^{-1}$. The density of the GCPC may be re-written as
@@
It is important to note that if the anisotropy parameter, $\lambda=1$, the GCPC distribution reduces to the circular independent projected Cauchy (CIPC) distribution (Tsagris et al., 2025)
@@
which is the WC distribution (Mardia and Jupp, 2000, pg. 51) with a different parameterization
@@
where $\gamma=\frac{2\delta}{1-\delta^{2}}$ or, conversely, $\delta=(\sqrt{\gamma^{2}+1}-1)/\gamma$ (Tsagris and Alzeley, 2025).
As the name suggests, $\lambda$ controls the anisotropy of the GCPC, and if $\lambda=1$ we end up with an isotropic covariance matrix $\bm{\Sigma}$.
@@ 2.1 Relationship with the CIPC distribution
The proof is straightforward by application of the change-of-variables formula. Without loss of generality, set $\omega=0$. Under the transformation $\psi=\arctan 2(\sin\varphi,\sqrt{\lambda}\cos\varphi)$ we can see that
$\cos\psi=\frac{\cos\varphi}{\sqrt{b}},\qquad\sin\psi=\frac{\sin\varphi}{\sqrt{\lambda b}},$
where $b=\cos^{2}\varphi+\sin^{2}\varphi/\lambda$. Differentiating $\tan\psi=\tan\varphi/\sqrt{\lambda}$ gives $|d\varphi/d\psi|=\sqrt{\lambda}\,b$. Applying the change-of-variables formula to $f_{\Phi}(\varphi)=(2\pi\sqrt{\lambda\,b}\,D)^{-1}$, with $D=c\sqrt{b}-\gamma\cos\varphi$ and $c=\sqrt{\gamma^{2}+1}$, we obtain:
$f_{\Psi}(\psi)\;=\;f_{\Phi}(\varphi)\cdot\sqrt{\lambda}b\;=\;\frac{\sqrt{b}}{2\pi D}\;=\;\frac{\sqrt{b}}{2\pi(c\sqrt{b}-\gamma\sqrt{b}\cos\psi)}\;=\;\frac{1}{2\pi(\sqrt{\gamma^{2}+1}-\gamma\cos\psi)},$
where in the last step we used $\cos\varphi=\sqrt{b}\cos\psi$. This is the $\mathrm{CIPC}(0,\gamma)\equiv\mathrm{WC}(0,\delta)$ density. ∎
Based on this we can define the opposite transformation as follows:
Lemma 2.1.
[…]
Proof.
The proof is again straightforward and hence omitted. ∎
2.2 Unimodality conditions for the GCPC distribution
Tsagris and Alzeley (2025) observed that the density function of the GCPC (8) may be bimodal. They equated the derivative of the log-density to zero and obtained:
$\sin(\theta-\omega)\left[2(\lambda-1)\sqrt{\gamma^{2}+1}\cos(\theta-\omega)\sqrt{\frac{(\lambda-1)\cos^{2}(\theta-\omega)+1}{\lambda}}-2(\lambda-1)\gamma\cos^{2}(\theta-\omega)-\gamma\right]$ $=$ $0.$ (19)
This yields two cases: either $\sin(\theta-\omega)=0$ or the expression in brackets equals zero. Case 1: $\sin(\theta-\omega)=0$.
$\theta-\omega=k\pi\implies\theta=\omega+k\pi,\ \ k\in\mathbb{Z}.$
Case 2: The term inside the brackets equals zero. Let $u=\cos(\theta-\omega)$:
$2(\lambda-1)\sqrt{\gamma^{2}+1}u\sqrt{\frac{(\lambda-1)u^{2}+1}{\lambda}}$ $=$ $\gamma\left(2(\lambda-1)u^{2}+1\right)$
$4(\lambda-1)^{2}(\gamma^{2}+1)\,u^{2}\frac{(\lambda-1)u^{2}+1}{\lambda}$ $=$ $\gamma^{2}\!\left(2(\lambda-1)u^{2}+1\right)^{2}\ \ \text{(By squaring both sides)}.$
Letting $t=u^{2}=\cos^{2}(\theta-\omega)$ and multiplying both sides with $\lambda$ yields:
$4(\lambda-1)^{3}(\gamma^{2}+1)\,t^{2}+4(\lambda-1)^{2}(\gamma^{2}+1)t$ $=$ $4\lambda\gamma^{2}(\lambda-1)^{2}t^{2}+4\lambda\gamma^{2}(\lambda-1)+\lambda\gamma^{2}.$
$4(\lambda-1)^{2}Mt^{2}+4(\lambda-1)Mt-\lambda\gamma^{2}$ $=$ $0,$
where $M=\lambda-\gamma^{2}-1$. Hence,
$t^{*}=\cos^{2}(\theta-\omega)=\frac{-M\pm\sqrt{M\left(M+\lambda\gamma^{2}\right)}}{2(\lambda-1)M}.$
Finally,
$\theta=\omega\pm\arccos\!\sqrt{t^{*}}.$
Notes.
  • If $M=0\implies\lambda=\gamma^{2}+1$, the quadratic degenerates and requires $\gamma=0$.
  • The term inside the square term of $t^{*}$ expands as $M\left(M+\lambda\gamma^{2}\right)=\left(\lambda-\gamma^{2}-1\right)\left[\left(\lambda-1\right)\left(\gamma^{2}+1\right)\right]$. To ensure that this is non-negative two conditions apply:
    • If $\lambda<1$ then $\lambda<\gamma^{2}+1$.
    • If $\lambda>1$ then $\lambda>\gamma^{2}+1$.
  • If $\lambda=1$, the term inside the brackets in (19) vanishes and we end up with Case 1 (unimodality of the distribution).
  • If $\lambda\neq 1$, then the following four cases apply:
    • Case A: $1<\lambda\leq\gamma^{2}+1$. The distribution is always unimodal since $M(\lambda-1)\leq 0$ gives complex roots, with no further conditions needed.
    • Case B: $\lambda>\gamma^{2}+1$. The distribution is unimodal when $t^{*}_{+}>1$, i.e. when:
      $(\gamma^{2}+1)(\lambda-1)>(\lambda-\gamma^{2}-1)(2\lambda-1)^{2}.$
    • Case C: $\frac{1}{2}<\lambda<1$. The distribution is unimodal when $t^{*}_{-}>1$, i.e. when:
      $(\gamma^{2}+1-\lambda)(1-2\lambda)^{2}>(\gamma^{2}+1)(1-\lambda).$
    • Case D: $\lambda\leq\frac{1}{2}$. Here $t^{*}_{-}\in[0,1]$ always, so the distribution is never unimodal for $\lambda\leq 1/2$ regardless of $\gamma$; bimodal critical points always exist.
The conditions stated in Cases A and D are straightforward for the bimodality. Cases B and C state that if $t^{*}$ falls outside the admissible region, the distribution is unimodal, where $t^{*}_{+}=\frac{M+\sqrt{M\left(M+\lambda\gamma^{2}\right)}}{2(\lambda-1)M}$ and $t^{*}_{-}=\frac{-M-\sqrt{M\left(M+\lambda\gamma^{2}\right)}}{2(\lambda-1)M}$. In those two cases, examination of bimodality requires some extra computations.
@@ 2.3 Probabilities
Following Tsagris and Alzeley (2025), instead of the cumulative probability function we derive the probability included within a given interval $(a,b)$, and provide an alternative formula.
@@ 2.4 Mean resultant length
The mean resultant length is defined as $\rho=E\left[\cos{\left(\theta-\omega\right)}\right]$. Based on this, one may compute the circular variance as $1-\rho$, and the circular standard deviation as $\left(-2\log{\rho}\right)^{1/2}$.
Theorem $2.$
The mean resultant length of the GCPC distribution is given by
$\rho=\frac{2(1+\delta)}{\pi(1-\delta)\sqrt{\lambda}}\;\Pi\left(\left(\frac{1+\delta}{1-\delta}\right)^{2}\frac{1}{\lambda}-1,1-\frac{1}{\lambda^{2}}\right),$
where $\Pi\left(n,k\right)$ is the complete elliptic integral of the third kind (Ward, 1960).
Proof.
$E[\cos\theta]=\frac{1}{2\pi\sqrt{\lambda}}\int_{-\pi}^{\pi}\frac{\cos\theta}{\sqrt{b}D}\,d\theta,\qquad b=\cos^{2}\theta+\frac{\sin^{2}\theta}{\lambda},\quad D=\sqrt{\gamma^{2}+1}\,\sqrt{b}-\gamma\cos\theta.$
Applying the change of variables $\psi=\arctan 2\!\left(\sin\theta,\,\sqrt{\lambda}\cos\theta\right)$ from Theorem 2.1, so that $f_{\mathrm{GCPC}}(\theta)\,d\theta=f_{WC}(\psi)\,d\psi$, we express $\cos\theta$ in terms of $\psi$:
$\cos\psi=\frac{\cos\theta}{\sqrt{b}},\qquad\sin\psi=\frac{\sin\theta}{\sqrt{\lambda b}},\qquad b=\frac{1}{1+(\lambda-1)\sin^{2}\psi},\qquad\cos\theta=\frac{\cos\psi}{\sqrt{1+(\lambda-1)\sin^{2}\psi}}.$
Substituting and using $f_{WC}(\psi)=\dfrac{1-\delta^{2}}{2\pi(1+\delta^{2}-2\delta\cos\psi)}$ we obtain:
$E\left(\cos\theta\right)=\frac{1-\delta^{2}}{2\pi}\int_{-\pi}^{\pi}\frac{\cos\psi}{\left(1+\delta^{2}-2\delta\cos\psi\right)\sqrt{1+(\lambda-1)\sin^{2}\psi}}\,d\psi.$
Since the integrand is even in $\psi$, doubling the integral over $0,\pi]$ and reducing via the substitution $u=\pi/2-\psi$ yields:
$E\left(\cos\theta\right)=\frac{2(1+\delta)}{\pi(1-\delta)\sqrt{\lambda}}\;\Pi\left(\left(\frac{1+\delta}{1-\delta}\right)^{\!2}\left(\frac{1}{\lambda}-1\right),\sqrt{1-\frac{1}{\lambda^{2}}}\right),$
confirming the non-analytical nature of $\rho=E[\cos\theta]$ for $\lambda\neq 1$. ∎
It remains the case that there is no analytical formula for $\rho$, unless $\lambda=1$, and this applies to higher trigonometric moments. Figure 1(a) visualizes the values of $\rho$ for a grid of values of $\gamma$ and $\lambda$. We observe that $\rho$ increases with increasing $\gamma$ and decreases with increasing $\lambda$ values.
@@ 2.5 Entropy
Note that if $\lambda=1$: $H_{GCPC}=\log(2\pi(1-\delta^{2}))=H_{WC}$, and if $\delta=0$: $H_{GCPC}=\log(8\pi\sqrt{\lambda}/(\sqrt{\lambda}+1)^{2})\leq\log(2\pi)$. Figure 1(b) visualizes the values of $H$ for a grid of values of $\gamma$ and $\lambda$. We observe that, in contrast to $\rho$, $H$ increases with increasing $\lambda$ and decreases with increasing $\gamma$ values.
@@ 2.6 Kullback-Leibler divergence between GCPC and CIPC
2.7 Maximum likelihood estimation
Tsagris and Alzeley (2025) performed maximum likelihood estimation (MLE) using the Euclidean representation of the density of the GCPC (7). We observed that this approach is sub-optimal and can yield a local instead of the global maximum, when the distribution is bimodal. We emphasize though that if the distribution is unimodal their approach is valid. Despite their efforts to ensure the maximum, via a clever implementation, we will show in the real data analysis, that the use of the log-likelihood parametrized using 7 does not lead to the global maximum, since the log-likelihood can have four local maxima [^1] Tsagris and Alzeley (2025) mentioned that the solution to Eq. (19) has four roots., when the distribution is bimodal. Our solution is to plot the data first and observe whether the distribution is unimodal or bimodal and then decide on which density parametrization to employ.
@@ 2.8.1 Hypothesis test for the location parameter of one sample
In order to test whether the location parameter equals some pre-specified value we will employ the log-likelihood ratio test. Under the $H_{0}$, $\omega=\omega_{0}$, and so we need to maximize the restricted log-likelihood, $\ell_{0}$, with respect to the other two parameters, $\gamma$ and $\lambda$, and estimate their values, $\tilde{\gamma}$ and $\tilde{\lambda}$. Under the $H_{1}$, $\omega\neq\omega_{0}$, and we maximize the unrestricted log-likelihood, $\ell_{1}$, to obtain $\hat{\omega}$, $\hat{\gamma}$ and $\hat{\lambda}$. Under regularity conditions, $\Lambda=2\left[\ell_{1}\left(\hat{\omega},\hat{\omega},\hat{\lambda}\right)-\ell_{0}\left(\omega_{0},\tilde{\gamma},\tilde{\lambda}\right)\right]\mathrel{\dot{\sim}}\chi^{2}_{1}$.
2.8.2 Confidence interval for the true location parameter
Using the log-likelihood ratio test, we can construct asymptotic confidence intervals for the true location parameter $\omega$ of a sample by searching for the pair of values for which it holds that (Cox and Hinkley, 1979): $\ell\left(\omega,\hat{\gamma},\hat{\lambda}\right)>\ell\left(\hat{\omega},\hat{\gamma},\hat{\lambda}\right)-\frac{\chi^{2}_{1}}{2}$.